# Scaling Efficient LLMs

### v5

@@ Abstract
Recent LLMs have hundreds of billions of parameters consuming vast resources. Furthermore, the so called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data. In response, we inquire into efficient LLMs, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus. Specifically, by comparing theoretical and empirical estimates of the Kullback-Leibler divergence, we derive a natural AI scaling law that the number of parameters in an efficient LLM scales as $D^{\gamma}$ where $D$ is the size of the training data and $\gamma\in[0.44,0.72]$, suggesting the existence of more efficient architectures. Against this backdrop, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Recurrent transformers (a) run in linear time in the sequence length, (b) are memory-efficient and amenable to parallel processing in large batches, (c) learn to forget history for language tasks, or accumulate history for long range tasks like copy and selective copy, and (d) are amenable to curriculum training to overcome vanishing gradients. In our experiments, we find that recurrent transformers perform favorably on benchmark tests.
@@ 1 Introduction
LLMs are artificial neural networks typically built on the transformer architecture as proposed by VSPUJGK17. Recent LLMs have hundreds of billions of parameters consuming vast amount of resources, e.g.,bmrskdoa20, rbcmhsoi21, spnlrcoc22, tdhskcol22. Furthermore, the so called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data, and that we are running out of usable training data to build even larger models, e.g., kmhbccgrwa20, hbmbcros22, muennighoff2023scaling. In response, we inquire into efficient LLMs independent of architecture, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus.
We compare theoretical and empirical estimates of the Kullback-Leibler divergence, kullback1951information, to show that at prevalent sequence lengths, the number of unique sequences in a natural training corpus of size $D$ is between $\Omega(D^{0.44})$ and $\mathcal{O}(D^{0.72})$. Our result leverages the empirical estimates of hbmbcros22. From a learning theoretic perspective, the number of unique sequences as a function of the size of a training corpus is an upper bound on the dimension of the space of functions required to learn any natural data set of that size. Theoretically, the number of parameters in an efficient LLM would scale linearly with the dimension of the space, and hence linearly with the number of unique sequences. In contrast to the so-called AI scaling law for the transformer architecture that suggests the number of parameters must scale linearly with the size of the data, a salient implication of our result is that the number of parameters in an efficient LLM scales as $D^{\gamma}$ for $0.44\leq\gamma\leq 0.72$, suggesting the existence of more efficient LLM architectures.
@@
With our theoretical results as backdrop, we seek more efficient LLM architectures. Preceding transformers, Recurrent Neural Networks (RNNs), e.g., elman1990finding, their variants Long Short-Term Memory (LSTM), e.g., hochreiter1997long and Gated Recurrent Units (GRUs), e.g., cho2014learning, were popular for linear time processing of sequential data streams such as natural language processing. More recent architectural alternatives include State Space Models, e.g., gu2021efficiently, and minimal RNNs, feng2024were. Building on prior work, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Prior work on recurrent transformers, e.g., dai2019transformer, where recurrence is used to extend context but does not participate in gradient calculations; and hutchins2022block, focuses on improving the long-range performance of multi-layer transformers by making some of the layers recurrent with large block sizes and many additional weights. In contrast, our focus is to find small and efficient networks for a given task, in the form of a single recurrent layer with a scalar parameter that learns to accumulate or forget history. Specifically, our recurrent transformers (a) run in linear time in the sequence length; (b) are memory-efficient and amenable to parallel processing in large batches (c) learn to forget history for language tasks, or accumulate history for long-range tasks like copy and selective copy; and (d) are amenable to curriculum training to overcome vanishing gradients, e.g.,hochreiter1991untersuchungen, sodhani2020toward. In our experiments, we find that a single-layer recurrent transformer can match the performance of multi-layer transformers on benchmark tests, delivering comparable accuracy with fewer parameters and much lower computational cost. Furthermore, the trainable accumulation capability enables a recurrent transformer to solve long range tasks at a fraction of the size of other recurrent architectures, e.g., gu2023mamba, ren2024exploring, and at a fraction of the cost of regular transformers.
@@ 2 Background
Let $S$ be the set of all sequences of length $l$ in the training corpus $T$. The Kullback-Leibler divergence of an LLM is
@@ 3 Empirical Scaling
Recent prior works report empirical expressions for the relationship between training loss and model scale, e.g., kmhbccgrwa20hbmbcros22, muennighoff2023scaling.  Per current practice, e.g., hbmbcros22, xfzzy23, these works train LLMs on a single epoch of natural training corpora.  Specifically, hbmbcros22 finds that for fixed sequence length, across a range of LLMs with $N$ parameters and training corpus of size $D$, the Kullback-Leibler divergence scales as follows
@@
where $\alpha=0.34,\beta=0.28,A=406.4$ and $B=410.7$ are fitting constants.
In the experiments, the training corpus is randomly split into a training subset and a test subset. The training subset is used for training the LLM, and the Kullback-Liebler divergence is computed on the test subset. Since each training run starts with a random initialization of the neural network, we assume Equation (3) holds with high confidence $(1-\delta)$, for small positive $\delta$ . The first term on the right reflects the ability of an LLM with $N$ parameters to capture the richness of the training corpus. The second term reflects the ability of the training corpus to capture the richness of the natural language.    Since $\alpha\approx\beta,A\approx B$, and the compute cost is $N\times D$, it follows that the optimal scaling strategy is $N\propto D$, which is also known as the AI Scaling Law.    However, $N\propto D$ suggests rote learning, raising the question of whether it is a natural requirement or a reflection of the transformer architecture used in the experiments.
@@ 4 Theoretical Scaling
We now consider LLMs in the context of the theoretical model of Probably Approximately Correct (PAC) learning, v84. Recall that $S$ is the set of sequences of length $l$ in $T$. Consider a hypothesis space of functions $F=\{f:S\rightarrow Y\}$, where for $s\in S$, $f(s)$ is a hypothesis $q_{s}$ for the probabilities $p_{s}$ of Equation 2. We follow sb14 in defining an agnostic learning algorithm as below. As in Section 3, the training corpus $T$ provides $D$ sequences of length $l$ as training samples for the learning algorithm. We define the loss function of the learning algorithm to be the Kullback-Leibler divergence between $f\in F$ and the natural distribution $p_{s}$ as follows
@@
Given $\epsilon,\delta\in(0,1)$ an agnostic learning algorithm finds $f\in F$ with confidence at least $(1-\delta)$ such that
$L(f,T)\leq\min_{h\in F}L(h,T)+\epsilon$
We assume that the natural distribution follows Zipf’s law, zipf2013psycho for natural languages, which is widely established and valid for sequences, ryland2015zipf.
Assumption 2.
The natural distribution follows Zips’s law in that $p_{s}\propto\frac{1}{r+Z}$, where $r$ is the rank of $s$ in descending order of frequency amongst $s\in S$, and $Z$ is a constant.
The following theorem uses the asymptotic complexity notations $\mathcal{O}$ and $\Omega$ for the upper and lower bound respectively.
@@
For any agnostic learning algorithm for $F$, $\epsilon$ is $\Omega(|S|/D)$ and $\mathcal{O}[(|S|/D)^{0.5}]$
@@
We first prove the upper bound. Let $\hat{p}_{s}$ and $\hat{L}(f,T)$ be the observed probabilities and loss respectively in a randomly chosen training subset of the corpus $T$. We have
$|\hat{L}(f,T)|=|\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/f(s))|\leq\sum_{s\in S}|\hat{p}_{s}log(\hat{p}_{s}/f(s))|=\sum_{s\in S}\hat{p}_{s}|log(\hat{p}_{s}/f(s))|$
Invoking Assumption 1 in the above, we get
$|\hat{L}(f,T)|\leq\sum_{s\in S}\hat{p}_{s}|log(\hat{p}_{s}/f(s))|\leq\sum_{s\in S}\hat{p}_{s}\lambda\leq\lambda$
@@
By Hoeffding’s inequality, feller1991introduction, we get
$\mathbb{P}\left\{\left|\hat{L}(f,T)-L(f,T)\right|>\epsilon\right\}\leq 2e^{-2D\epsilon^{2}/\lambda}$
Summing over all choices for $f$, we get
$\sum_{f\in F}\mathbb{P}\left\{\left|\hat{L}(f,T)-L(f,t)\right|>\epsilon\right\}\leq|Y|^{|S|}2e^{-2D\epsilon^{2}/\lambda}$
We wish to bound $\delta$ with the RHS above, i.e., $\delta\leq|Y|^{|S|}2e^{-2D\epsilon^{2}/\lambda}$. Taking logarithms on both sides and treating $\delta,\lambda$, and $|Y|$ as constants yields $\epsilon=\mathcal{O}[(|S|/D])^{0.5}$.
Next we prove the lower bound. For any $f\in F$,
$L(f,T)-\hat{L}(f,T)$ $=\sum_{s\in S}p_{s}log(p_{s}/f(s))-\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/f(s))$
$=\sum_{s\in S}p_{s}log(p_{s}/\hat{p}_{s})-\sum_{s\in S}(\hat{p}_{s}-p_{s})log(\hat{p}_{s})+\sum_{s\in S}(\hat{p}_{s}-p_{s})log(f(s))$
Taking the expectation $\mathbb{E}$ on both sides of the above with respect to random training subsets of T, and noting that $\mathbb{E}\{(\hat{p}_{s}-p_{s})\}=0$, we get,
$\mathbb{E}\left\{L(f,T)-\hat{L}(f,T)\right\}=\mathbb{E}\left\{\sum_{s\in S}p_{s}log(p_{s}/\hat{p}_{s})\right\}$ (4)
Since each sequence $s$ is a Bernoulli variable in a random training subset, the variance of $\hat{p}_{s}$ is $p_{s}(1-p_{s})$. Therefore by the Central Limit Theorem, e.g., feller1991introduction, as $D$ increases without bound, almost surely,
$|\hat{p}_{s}-p_{s}|\propto\sqrt{\frac{p_{s}(1-p_{s})}{D}}$
Invoking Assumption 2 in the above and summing over $s\in S$, we get
$\sum_{s\in S}|\hat{p}_{s}-p_{s}|\propto\frac{1}{\sqrt{D}}\sum_{r=1}^{|S|}\sqrt{\frac{1-\frac{1}{r+Z}}{r+Z}}$
Approximating the sum in the right-hand-side above with the definite integral, we get that
$\sum_{s\in S}|(\hat{p}_{s}-p_{s}|\propto\sqrt{\frac{|S|}{D}}$
Combining the above with Pinsker’s inequality, csiszar2011information, we get
$\sum_{s\in S}p_{s}log(p_{s}/\hat{p}_{s})\geq\frac{1}{2}\left(\sum_{s\in S}|(\hat{p}_{s}-p_{s}|\right)^{2}=\Omega[|S|/D]$
Combining the above with Equation 4, we set
$\zeta=\mathbb{E}\left\{L(f,T)-\hat{L}(f,T)\right\}=\Omega[|S|/D]$
By Assumption 1, $L(f,T)-\hat{L}(f,T)\leq 2\lambda$. Let $x$ be the probability that $L(f,T)-\hat{L}(f,T)>\zeta/2$. The minimum value of $x$ is achieved by placing a probability mass of $x$ at $2\lambda$ and (1x)(1-x) at $\zeta/2$. Hence
$2\lambda x+(\zeta/2)(1-x)=\zeta\implies x(2\lambda-\zeta/2)=\zeta/2\implies x=\frac{\zeta}{4\lambda-\zeta}>\frac{\zeta}{4\lambda}$
It follows that with probability greater than $\frac{\zeta}{4\lambda}$,
$L(f,T)-\hat{L}(f,T)>\zeta/2=\Omega[|S|/D]$
implying that for $f\in F$ output by a learning algorithm, $L(f,T)>\min_{h\in F}L(h,T)+\Omega[|S|/D]$. Hence the theorem.
@@
We now link the empirical bounds of Equation 3 with the theorem above to estimate the number of unique sequences in a training corpus as a function of its size. Our analysis compares the empirical estimates of hbmbcros22 with the estimates of Theorem 1 to extract properties of the underlying natural language.
@@
If Equation 3 holds for sequences of length $l$, the number of unique sequences $|S|$ of length $l$ in a corpus of size $D$ is between $\Omega[D^{0.44}|]$ and $\mathcal{O}[D^{0.72}]$
@@
Consider an LLM with $N$ parameters representing a space of functions $F=\{f:S\rightarrow Y\}$. Suppose the LLM computes a function $f$ after training on corpus $T$.  From Equation 2 and the upper bound of Theorem 1, we have
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/q_{s})=\min_{h\in F}L(h,T)+\mathcal{O}[(|S|/D)^{0.5}]$
@@
$\Delta=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}=\min_{h\in F}L(h,T)+\mathcal{O}[(|S|/D)^{0.5}]$
As the number of parameters $N\rightarrow\infty$, $A/N\rightarrow 0$, and since transformers are universal approximators, $\min_{h\in F}L(h,T)\rightarrow 0$. Hence we have
$\frac{B}{D^{\beta}}=\mathcal{O}[(|S|/D)^{0.5}]$
Substituting $\beta=0.28$ from Equation 3 in the above,  we get
@@
Next, we combine Equations 2 with the lower bound of Theorem 1.
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/q_{s})=\min_{h\in F}L(h,T)+\Omega[|S|/D]$
@@
$\Delta=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}=\min_{h\in F}L(h,T)+\Omega[|S|/D]$
As the number of parameters $N\rightarrow\infty$, $A/N\rightarrow 0$, and since transformers are universal approximators, $\min_{h\in F}L(h,T)\rightarrow 0$. Hence we have
$\frac{B}{D^{\beta}}=\Omega[|S|/D]$
@@
Which completes the proof.
From a learning theoretic perspective, the number of unique sequences as a function of the size of a training corpus is an upper bound on the degrees of freedom required to learn any natural dataset of that size. Theoretically, the number of parameters in an efficient LLM would scale linearly in the number of unique sequences; e.g., a "brute-force" code book of one entry per unique sequence. The so-called AI scaling law for the transformer architecture suggests that the number of parameters must scale linearly with the size of the data, e.g., kmhbccgrwa20, hbmbcros22, muennighoff2023scaling. In contrast, our main result implies that the number of parameters of an efficient LLM need only scale as $D^{\gamma}$ for $\gamma\in[0.44,0.72]$, suggesting the existence of substantially more efficient LLM architectures.

### v6

@@ Abstract
Recent LLMs have hundreds of billions of parameters consuming vast resources. Furthermore, the so-called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the length of the data. In response, we inquire into efficient LLMs, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus. Specifically, by comparing theoretical and empirical estimates of the Kullback-Leibler divergence, we derive a natural AI scaling law that the number of parameters in an efficient LLM scales as $D^{\gamma},\gamma\in[0.44,0.72]$ where $D$ is the length of the training data, suggesting the existence of more efficient architectures. Against this backdrop, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Recurrent transformers (a) run in linear time in the sequence length, (b) are memory-efficient and amenable to parallel processing in large batches, (c) learn to forget history for language tasks, or accumulate history for long range tasks like copy and selective copy, and (d) are amenable to curriculum training to overcome vanishing gradients. In our experiments, we find that recurrent transformers perform favorably on benchmark tests.
@@ 1 Introduction
LLMs are artificial neural networks typically built on the transformer architecture as proposed by VSPUJGK17. Recent LLMs have hundreds of billions of parameters consuming vast amount of resources, e.g.,bmrskdoa20, rbcmhsoi21, spnlrcoc22, tdhskcol22. Furthermore, the so-called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the length of the data, and that we are running out of usable training data to build even larger models, e.g., kmhbccgrwa20, hbmbcros22, muennighoff2023scaling. In response, we inquire into efficient LLMs independent of architecture, i.e., those with the fewest parameters that achieve the desired accuracy on a training corpus.
Our method of inquiry is via estimating the learning theoretic dimensionality of the space of models required for a language, which is bounded by the number of unique sequences of at most a given length in a training corpus, as a function of its length. Since direct empirical estimates of such requires vast resources, our analysis takes an alternate approach, comparing theoretical estimates of entropic loss with published empirical estimates. Specifically, we propose a learning theoretic framework for LLMs, in the context of Probably Approximately Correct (PAC) learning, v84; sb14. In this framework, we compare theoretical and empirical estimates of the entropic loss, to show that at prevalent sequence lengths, the number of unique sequences in a natural training corpus of length $D$ scales approximately as $D^{\gamma},\gamma\in[0.44,0.72]$. Our result leverages the empirical estimates of hbmbcros22. From a learning theoretic perspective, the number of unique sequences as a function of the length of a training corpus is an asymptotic upper bound on the dimension of the space of functions required to learn any natural data set of that length. Theoretically, the number of parameters in an efficient LLM would scale linearly with the dimension of the space, and hence linearly with the number of unique sequences. In contrast to the so-called AI scaling law for the transformer architecture that suggests the number of parameters must scale linearly with the length of the data, a salient implication of our result is that the number of parameters in an efficient LLM scales as $D^{0.44}$, suggesting the existence of more efficient LLM architectures.
@@
With our theoretical results as backdrop, we seek more efficient LLM architectures. Preceding transformers, Recurrent Neural Networks (RNNs), e.g., elman1990finding, their variants Long Short-Term Memory (LSTM), e.g., hochreiter1997long and Gated Recurrent Units (GRUs), e.g., cho2014learning, were popular for linear time processing of sequential data streams such as natural language processing. More recent architectural alternatives include State Space Models, e.g., gu2021efficiently, and minimal RNNs, feng2024were. Building on prior work, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Prior work on recurrent transformers, e.g., dai2019transformer, where recurrence is used to extend context but does not participate in gradient calculations; and hutchins2022block, focuses on improving the long-range performance of multi-layer transformers by making some of the layers recurrent with large block sizes and many additional weights. In contrast, we focus on finding small and efficient networks for a given task, in the form of a single recurrent layer with a scalar parameter that learns to accumulate or forget history. Specifically, our recurrent transformers (a) run in linear time in the sequence length; (b) are memory-efficient and amenable to parallel processing in large batches (c) learn to forget history for language tasks, or accumulate history for long-range tasks like copy and selective copy; and (d) are amenable to curriculum training to overcome vanishing gradients, e.g.,hochreiter1991untersuchungen, sodhani2020toward. In our experiments, we find that a single-layer recurrent transformer can match the performance of multi-layer transformers on benchmark tests, delivering comparable accuracy with fewer parameters and much lower computational cost. Furthermore, the trainable accumulation capability enables a recurrent transformer to solve long range tasks at a fraction of the size of other recurrent architectures, e.g., gu2023mamba, ren2024exploring, and at a fraction of the cost of regular transformers.
@@ 2 Background
Let $S$ be the set of unique sequences of length at most $l$ in $T$. The Kullback-Leibler divergence, kullback1951information, of an LLM is
@@ 3 Empirical Scaling
Recent prior works report empirical expressions for the relationship between training loss and model scale, e.g., kmhbccgrwa20hbmbcros22, muennighoff2023scaling.  Per current practice, e.g., hbmbcros22, xfzzy23, these works train large LLMs that overfit for a single epoch of natural training corpora.  Specifically, hbmbcros22 finds that for a sequence length of 2048 tokens, across hundreds of LLMs and corpus sizes, the Kullback-Leibler divergence scales as follows
@@
where $N$ is the number of parameters, $D$ the corpus length, and $\alpha=0.34,\beta=0.28,A=406.4$ and $B=410.7$ are fitting constants.
In the experiments, the given corpus is randomly split into a training subset and a test subset of fixed proportions. The training subset is used for training the LLM, and the cross-entropy loss is computed on the test subset. Each training run starts with a random initialization of the neural network, and Equation (3) represents the high-confidence outcome. The first term on the right is the inherent entropy of the language. The second term on the right reflects the ability of an LLM with $N$ parameters to capture the richness of the training corpus. The third term reflects the ability of the training corpus to capture the richness of the natural language.    Since $\alpha\approx\beta,A\approx B$, and the compute cost is $N\times D$, it follows that the optimal scaling strategy is $N\propto D$, which is also known as the AI Scaling Law.    However, $N\propto D$ suggests rote learning, raising the question of whether it is a natural requirement or a reflection of the transformer architecture used in the experiments.
@@ 4 Theoretical Scaling
We now extend the framework of Probably Approximately Correct (PAC) learning, v84; sb14 to LLMs through the following definitions across a given alphabet and sequence length $l$.
Definition 1.
Let $S$ be the set of sequences of length at most $l$ on the alphabet. A Large Language Model (LLM) is a probability distribution $f:S\rightarrow Y$. A class of LLMs is the set of all such distributions.
In Section 3, a corpus is loosely described as a continuous stream of text. Formally, we define a corpus as a collections of sequences of length at most $l$ tokens. Furthermore, during training, the corpus is said to be split into a "training subset" and "test subset" of fixed proportions. Formally, the "training subset" and "test subset" are actually randomly chosen sub-collections of the corpus.
Definition 2.
A corpus $T$ is a collection of sequences of length at most $l$. A training subset of $T$ is a randomly chosen sub-collection that includes any sequence in $T$ with a specified sampling probability $t$.
Definition 3.
A learning algorithm for a class of LLMs $F$, takes as input a training subset of a corpus $T$ and produces as output $f\in F$, such that $f(s)$ is a hypothesis $q_{s}$ for the natural probabilities $p_{s}$, as in Equation 2. The algorithm seeks to minimize the Kullback-Leibler divergence between $f$ and the natural distribution $p_{s}$ across the set $S$ of unique sequences in $T$ as follows
@@
The excess risk of the learning algorithm is $L(f,T)-\min_{h\in F}L(h,T)$.
Recall from Section 3 that in practice LLMs are over sized and overfit the training data in that the training loss is less than the test loss. The following definition formally characterizes overfitting.
@@
[…]
The following theorem uses the asymptotic complexity notations $\mathcal{O}$ and $\Omega$ for the upper and lower bound respectively. The theorem is stated in terms of the number of unique sequences in a given corpus, which is an upper bound on the Natarajan dimension, n89; sb14 of a class of LLMs on the corpus from a learning theoretic perspective, and the number of degrees of freedom, akaike1998information, from an information theoretic perspective.
@@
Given are a sequence length $l$, a class of LLMs $F$, and a corpus $T$. Let $S$ be the set of sequences in $T$. (a) There exists a learning algorithm for $F$, such that for a given $\delta>0$, with probability $(1-\delta)$ the excess risk is $\mathcal{O}\left[\sqrt{\frac{|S|}{|T|}}\right]$; (b) If $F$ overfits $T$, then for any learning algorithm for $F$, for a given $\delta>0$, with probability $(1-\delta)$, the excess risk is $\Omega\left[\frac{|S|}{|T|}\right]-\mathcal{O}\left[\frac{1}{\sqrt{|T|}}\right]$.
@@
[…]
$=\left(\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s})-\sum_{s\in S}p_{s}log(p_{s})\right)-\left(\sum_{s\in S}\hat{p}_{s}log(f(s))-\sum_{s\in S}p_{s}log(f(s))\right)$
Which implies
$|L(f,\hat{T})-L(f,T)|$ $\leq\left|\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s})-\sum_{s\in S}p_{s}log(p_{s})\right|$
$+\left|\sum_{s\in S}\hat{p}_{s}log(f(s))-\sum_{s\in S}p_{s}log(f(s))\right|$ (5)
Now
$\sum_{s\in S}\hat{p}_{s}log(f(s))=\sum_{s\in\hat{T}}\hat{p}_{s}log(f(s))=\sum_{s\in\hat{T}}\frac{1}{|\hat{T}|}log(f(s)),$
and taking the expectation over all possible random training subsets $\hat{T}$ of $T$,
$\mathbb{E}\left\{\sum_{s\in\hat{T}}\frac{1}{|\hat{T}|}log(f(s))\right\}=\sum_{s\in S}p_{s}log(f(s)).$
@@
By Assumption 1, $|log(f(s))|\leq\lambda$ in the above, and hence by Hoeffding’s inequality, feller1991introduction,
$\mathbb{P}\left\{\left|\sum_{s\in\hat{T}}\frac{1}{|\hat{T}|}log(f(s))-\sum_{s\in S}p_{s}log(f(s))\right|>\zeta\right\}\leq 2e^{-2|\hat{T}|\zeta^{2}/\lambda^{2}}$
Since $F:S\rightarrow Y$, $|F|\leq|Y|^{|S|}$. Summing over all choices for $f$, we get
$\sum_{f\in F}\mathbb{P}\left\{\left|\sum_{s\in\hat{T}}\frac{1}{|\hat{T}|}log(f(s))-\sum_{s\in S}p_{s}log(f(s))\right|>\zeta\right\}\leq|Y|^{|S|}2e^{-2|\hat{T}|\zeta^{2}/\lambda^{2}}$
Setting $\delta\leq|Y|^{|S|}2e^{-2|\hat{T}|\zeta^{2}/\lambda^{2}}$, noting $|\hat{T}|=t|T|$, and treating $t,\delta,\lambda$, and $|Y|$ as constants, we get that with probability at least $(1-\delta),\;\zeta=\mathcal{O}\left[\sqrt{\frac{|S|}{|T|}}\right]$.
Let $H=-\sum_{s\in S}p_{s}log(p_{s})$ denote the entropy of the natural distribution. By Assumption 1, $|log(p_{s})|\leq\lambda$ in the above, and hence the variance $Var(H)\leq\lambda^{2}$. By the Central Limit Theorem, feller1991introduction, for given $\delta$, with probability $(1-\delta)$,
$\eta=\left|\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s})-\sum_{s\in S}p_{s}log(p_{s})\right|=\mathcal{O}\left[\sqrt{\frac{Var(H)}{|\hat{T}|}}\right$]
Combining the above while noting $|\hat{T}|\approx t|T|$, and treating $t,\delta$ and $\lambda$ as constants yields $\eta=\mathcal{O}\left[\sqrt{\frac{1}{|T|}}\right]$.
Returning to Equation 4, we have that with probability at least $(1-\delta)$
$|L(f,\hat{T})-L(f,T)|\leq\eta+\zeta=\mathcal{O}\left[\sqrt{\frac{|S|}{|T|}}\right]$
[…]
$L(g,\hat{T})-L(f,T)\leq L(f,\hat{T})-L(f,T)=\mathcal{O}\left[\sqrt{\frac{|S|}{|T|}}\right]$
completing the proof of Part (a).
[…]
Subtracting $L(f,T)\leq L(g,T)$ from the above, we get
$L(g,T)-L(f,T)$ $\geq L(g,\hat{T})-L(g,T)$
$\geq\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/g(s))-\sum_{s\in S}p_{s}log(p_{s}/g(s))$
Adding and subtracting $\sum_{s\in S}\hat{p}_{s}log(p_{s})$ from the right-hand side,
$L(g,T)$ $-L(f,T)$
$\geq\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/g(s))-\sum_{s\in S}\hat{p}_{s}log(p_{s})-\sum_{s\in S}p_{s}log(p_{s}/g(s))+\sum_{s\in S}\hat{p}_{s}log(p_{s})$
$\geq\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/p_{s})+\sum_{s\in S}(\hat{p}_{s}-p_{s})log(p_{s}/g(s))$
Taking expectations on both sides over the empirical distribution $\hat{p}_{s}$, we get
$\mathbb{E}\left\{L(g,T)-L(f,T)\right\}$ $\geq\mathbb{E}\left\{\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/p_{s})\right\}\approx\frac{|S|-1}{2|\hat{T}|}=\Omega\left[\frac{|S|}{|T|}\right]$ (6)
The approximation above, kullback1997information, follows the "delta method" of a Taylor series in two terms for $log(\hat{p}_{s}/p_{s})$ and thereafter substitutes the variance of the multinomial distribution in the expectation.
By construction and Assumption 1, $L(g,T)-L(f,T)\in[0,\lambda]$ and hence $\mathbb{Var}\{L(g,T)-L(f,T)\}\leq\lambda^{2}$. Therefore, combining Equation 6 with the Central Limit Theorem, for given $\delta$, with probability $(1-\delta)$,
$\left|(L(g,T)-L(f,T))-\mathbb{E}\left\{L(g,T)-L(f,T)\right\}\right|$ $\leq\mathcal{O}\left[\frac{\lambda}{\sqrt{\hat{|T|}}}\right$]
Treating $\lambda$ as a constant, we get that for given $\delta$, with probability $(1-\delta)$,
$L(g,T)-L(f,T))\geq\mathbb{E}\left\{L(g,T)-L(f,T)\right\}-\mathcal{O}\left[\frac{1}{\sqrt{|T|}}\right]$
Substituting Equation 6, we get that for given $\delta$, with probability $(1-\delta)$,
$L(g,T)-L(f,T))=\Omega\left[\frac{|S|}{|T|}\right]-\mathcal{O}\left[\frac{1}{\sqrt{|T|}}\right]$ (7)
This completes the proof. ∎
@@
We now link the empirical bounds of Equation 4 with the theorem above to estimate the number of unique sequences in a training corpus as a function of its length. Our analysis combines the empirical estimates of hbmbcros22 with the estimates of Theorem 1 to extract properties of the underlying natural language.
@@
If Equation 4 holds for sequence length $l$, the number of unique sequences of length at most $l$ in a corpus of length $D$ is between $\Omega[D^{0.44}]$ and $\mathcal{O}[D^{0.72}]$.
@@
Consider a transformer with $N$ parameters as in Section 3, computing a class of functions $F:S\rightarrow Y$. Suppose that after training on a corpus of length $D$ yielding a collection $T$ of $lD$ sequences of length at most $l$ tokens, the transformer computes a distribution $f\in F$.  From Equation 2 and the upper bound of Theorem 1 and treating $l$ as a constant, we have
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/f(s))=\min_{h\in F}L(h,T)+\mathcal{O}\left[\sqrt{\frac{|S|}{|T|}}\right]=\min_{h\in F}L(h,T)+\mathcal{O}\left[\sqrt{\frac{|S|}{D}}\right]$
@@
$\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}=\min_{h\in F}L(h,T)+\mathcal{O}\left[\sqrt{\frac{|S|}{D}}\right]$
As the number of parameters $N\rightarrow\infty$, $A/N\rightarrow 0$, and since transformers are universal approximators, $\min_{h\in F}L(h,T)\rightarrow 0$, yielding
$\frac{B}{D^{\beta}}=\mathcal{O}\left[\sqrt{\frac{|S|}{D}}\right]$
Substituting $\beta=0.28$ from Section 3 in the above,  we get
@@
Next, we note that the LLMs of Section 3 overfit and hence satisy the condition of Theorem 1. Furthermore, we assume that $|S|$ is asymptotically larger than $\sqrt{D}$, implying that $\Omega\left[\frac{|S|}{|T|}\right]$ is the dominant term in the lower bound of Theorem 1. We verify that the assumption is not contradicted when we combine Equations 2 with the lower bound of Theorem 1 as below.
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/f(s))=\min_{h\in F}L(h,T)+\Omega\left[\frac{|S|}{|T|}\right]=\min_{h\in F}L(h,T)+\Omega\left[\frac{|S|}{D}\right]$
@@
$\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}=\min_{h\in F}L(h,T)+\Omega\left[\frac{|S|}{D}\right]$
As the number of parameters $N\rightarrow\infty$, $A/N\rightarrow 0$, and since transformers are universal approximators, $\min_{h\in F}L(h,T)\rightarrow 0$, yielding
$\frac{B}{D^{\beta}}=\Omega\left[\frac{|S|}{D}\right]$
@@
The above does not contradict our assumption that $|S|$ is asymptotically larger than $\sqrt{D}$, completing the proof.
From a learning theoretic perspective, the number of unique sequences as a function of the length of a training corpus is an upper bound on the degrees of freedom required to learn any natural dataset of that length. Theoretically, the number of parameters in an efficient LLM would scale linearly in the number of unique sequences; e.g., a "brute-force" code book of one entry per unique sequence. The so-called AI scaling law for the transformer architecture suggests that the number of parameters must scale linearly with the length of the data, e.g., kmhbccgrwa20, hbmbcros22, muennighoff2023scaling. In contrast, our main result implies that the number of parameters of an efficient LLM need only scale as $D^{\gamma},\gamma\in[0.44,0.72]$, suggesting the existence of substantially more efficient LLM architectures.